OpenStudy (anonymous):

Solve t = p+3r over 6 for r.

OpenStudy (anonymous):

Keep It Simple Stupid Kiss The part of the question making it look the hardest is the far cation so remove is by multiplying both sides by 6

OpenStudy (anonymous):

Then you have 6t= p+3r

OpenStudy (anonymous):

There are two terms on the right hand side, one with the r and one without an r, so remove the one without the r and put it on the other side Minus p

OpenStudy (anonymous):

so 6t - p = 3r

OpenStudy (anonymous):

Then you have the 3r, to get r on its own you divide by 3

OpenStudy (anonymous):

so \[\frac{ 6t-p }{ 3 } = r\]

OpenStudy (anonymous):

this is assuming the original question was \[t=\frac{ p+3r }{ 6 }\]

OpenStudy (anonymous):

Hope it makes sense :)

OpenStudy (anonymous):

yes it does thank you!